# May 2, 2024

### v1

@@ 4.1.2 The Heisenberg group deformation
which is isomorphic to the Heisenberg algebra. Following a similar procedure as before, we start with the faithful representation
@@
The compatible Poisson-Lie structure can be computed directly, and its fundamental Poisson brackets read
$\begin{array}[]{lll}\{x,y\}_{\eta}=0,&\{x,z\}_{\eta}=\eta\,x+y&\{y,z\}_{\eta}=x+\eta\,y.\end{array}$ (4.13)
@@
Again, it is straightforward to check that in the limit $\eta\to 0$ we recover the Poisson version of the $(1+1)$-dimensional Poincaré Lie algebra (3.5) and the Casimir (2.6).
@@ 4.2 Bi-Hamiltonian deformations
Integrable deformations of Case B coming independently from each of its two Lie-Poisson Hamiltonian structures could be obtained by a completely analogous procedure to the one presented above, and will be based on the complete classification of Lie bialgebra structures for the centrally extended (1+1) Poincaré and $so(3)$ algebras. For the sake of brevity we shall not discuss them in this paper in detail. However, Rikitake system B is bi-Hamiltonian and thus the existence of the Poisson pencil (2.18) allows in principle to construct bi-Hamiltonian integrable deformations by following the prodedure introduced in [7, 8]. The essential point is that these deformations exist if and only if there are common cocycles $\delta$ for the two Lie algebras underlying the bi-Hamiltonian structure, namely the centrally extended Poincaré and $so(3)$ Lie algebras. In the following we prove that Case B do not admit bi-Hamiltonian integrable deformations since this condition cannot be fulfilled. Nevertheless, later we show how Case AB does admit such a common cocycle, and we compute and analyse a particular bi-Hamiltonian deformation for this system.
@@
We set a generic map $\delta:\mathfrak{g}\to\mathfrak{g}\wedge\mathfrak{g}$
$\begin{array}[]{l}\delta(X)=\eta\,(a_{1}\,X\wedge Y+a_{2}\,X\wedge Z+a_{3}\,X\wedge W+a_{4}\,Y\wedge Z+a_{5}\,Y\wedge W+a_{6}\,Z\wedge W)\\ \delta(Y)=\eta\,(b_{1}\,X\wedge Y+b_{2}\,X\wedge Z+b_{3}\,X\wedge W+b_{4}\,Y\wedge Z+b_{5}\,Y\wedge W+b_{6}\,Z\wedge W)\\ \delta(Z)=\eta\,(c_{1}\,X\wedge Y+c_{2}\,X\wedge Z+c_{3}\,X\wedge W+c_{4}\,Y\wedge Z+c_{5}\,Y\wedge W+c_{6}\,Z\wedge W)\\ \delta(W)=\eta\,(d_{1}\,X\wedge Y+d_{2}\,X\wedge Z+d_{3}\,X\wedge W+d_{4}\,Y\wedge Z+d_{5}\,Y\wedge W+d_{6}\,Z\wedge W)\\ \end{array}$ (4.17)
and from here we impose the co-cycle condition
@@
$\begin{array}[]{l}\delta(X)=\eta\left((1-2\lambda)b_{3}\,Y\wedge W+b_{4}\,X\wedge Z-b_{4}\dfrac{\beta(1-\lambda)}{\lambda}X\wedge W+(2\beta(1-\lambda)c_{2}-2\lambda c_{3})Z\wedge W-c_{4}\,X\wedge Y\right)\\ \\ \delta(Y)=\eta\left(b_{3}\,X\wedge W+b_{4}\,Y\wedge Z-b_{4}\dfrac{\beta(\lambda-1)}{\lambda}Y\wedge W+b_{6}Z\wedge W+c_{2}X\wedge Y\right)\\ \\ \delta(Z)=\eta\left(c_{2}X\wedge Z+c_{3}X\wedge W+c_{4}Y\wedge Z+\left(\dfrac{(2\lambda-1)}{2\lambda}b_{6}+\dfrac{\beta(1-\lambda)}{\lambda}c_{4}\right)Y\wedge W\right)\\ \\ \delta(W)=0.\end{array}$ (4.19)
Since $\delta$ should be defined for the complete Poisson pencil, including $\lambda=0$, we obtain that $b_{4}=0$, $b_{6}=0$, and $c_{4}=0$. Therefore we have
$\begin{array}[]{l}\delta(X)=\eta\left((1-2\lambda)b_{3}\,Y\wedge W+(2\beta(1-\lambda)c_{2}-2\lambda c_{3})Z\wedge W\right)\\ \\ \delta(Y)=\eta\left(b_{3}\,X\wedge W+c_{2}X\wedge Y\right)\\ \\ \delta(Z)=\eta\left(c_{2}X\wedge Z+c_{3}X\wedge W\right)\\ \\ \delta(W)=0.\end{array}$ (4.20)
The only remaining $\lambda$-independent solution is obtained if $b_{3}=0$ and $c_{3}=-\beta c_{2}$,
@@
However, the co-Jacobi condition implies that the only cocommutator map is the trivial one $\delta(X)=\delta(Y)=\delta(Z)=\delta(W)=0$. Therefore, there are no non-trivial bi-Hamiltonian integrable deformations of the system (4.15).
@@ 4.2.1 Bi-Hamiltonian deformation of the Rikitake AB system
We have seen that Case A does not admit (non-trivial) bi-Hamiltonian deformations. However, when $\alpha=\mu=\beta=0$ (Case AB), the system (2.19) does admit such deformations. This case is given by the system of ODEs
@@
    First Hamiltonian structure: The (1+1)-Poincaré Lie algebra
    $\begin{array}[]{lll}\{x,y\}_{0}=0,&\{x,z\}_{0}=y,&\{y,z\}_{0}=x\end{array}$ (4.23)
    where the Casimir is given by $\mathcal{C}_{0}=\dfrac{1}{4}(y^{2}-x^{2})$, together with the Hamiltonian function
@@
    Second Hamiltonian structure: The $\mathfrak{so}(3)$ Lie algebra
    $\begin{array}[]{lll}\{x,y\}_{1}=2z,&\{x,z\}_{1}=-y,&\{y,z\}_{1}=x\end{array}$ (4.25)
@@
It is easy to see that, in contradistinction to Case A, now there are $\lambda$-independent common cocycles, like for instance, the cocycle given in (3.6),
@@
The ‘book’ Lie group can be embedded in $\mathrm{GL}(3,\mathbb{R})$ as shown in (4.2). Using this parametrisation, the coproduct $\Delta_{\eta}:\mathcal{C}^{\infty}(G^{\ast})\to\mathcal{C}^{\infty}(G^{\ast})\otimes\mathcal{C}^{\infty}(G^{\ast})$ is given by (4.4). A standard computation shows that the unique Poisson structure on $G^{\ast}$ compatible with $\Delta_{\eta}$ and therefore endowing $\mathcal{C}^{\infty}(G^{\ast})$ with a Poisson-Hopf algebra is given by the fundamental Poisson brackets
@@
It is easy to check that the Casimir function for this Poisson-Lie group structure is given by
@@
When $\lambda=0$ and $\lambda=1$, these structures are Poisson-Lie deformations of the Lie-Poisson (1+1)-Poincaré (4.23) and $\mathfrak{so}(3)$ (4.25) Lie algebras, respectively.
@@
    As it was expected the only deformation in the Poisson-Lie brackets is produced for the $\{x,y\}_{0,\eta}$ bracket whose $\eta\rightarrow 0$ limit trivially gives us back the $\{x,y\}_{0,0}=0$. In this case the deformed Casimir (4.31) reduces to
@@
    Moreover, the Casimir function reads
@@ 5.1 Case A
$\mathcal{H}=\dfrac{1}{2}(x^{2}+z^{2})-\alpha\,\log(x+y).$ (5.1)
The undeformed system (2.5) is obtained from the (1+1) Poincaré algebra (2.4), as shown in Section 2. This system could be seen as a dynamical system defined on an Abelian (trivial cocommutator) Poisson-Lie group. Therefore $\eta$-deformations, or equivalently, non-Abelian dual Poisson-Lie structures, define integrable deformations of this system. In this case, both Poisson-Lie structures explicitly given in the previous Section define the following deformed dynamical systems.
The “book” group deformation:
@@
A symplectic realisation of this deformed system is given by
@@
Note that in the $\eta\to 0$ limit we recover (2.7), both in the dynamical system and the symplectic realisation.
Refer to captionRefer to caption
@@
The Heisenberg group deformation:
@@
$\begin{split}\dot{x}&=\{x,\mathcal{H}\}_{\eta}=z\,(y+\eta\,x)\\ \dot{y}&=\{y,\mathcal{H}\}_{\eta}=z\,(x+\eta\,y)\\ \dot{z}&=\{z,\mathcal{H}\}_{\eta}=\alpha-x\,y+\eta(\alpha-x^{2}).\end{split}$ (5.4)
A posible symplectic realisation in this case is given by
@@
The most significant difference between (5.3) and (5.5) is that in the former both $x$ and $y$ are functions of $q$ and $p$, while in the latter they are only functions of the $q$ variables.
Refer to captionRefer to caption
@@
Note that trajectories of the system with large positive values of the deformation parameter $\eta$ tend to be contained in a plane, while for large negative values of the deformation parameter $\eta$ they bend themselves and are truly 3-dimensional.
@@ 5.2 Case AB
In order to compute the deformed Hamilton’s equations we have to take as Hamiltonian
$\mathcal{H}_{1,\eta}=\mathcal{C}_{0,\eta}=\dfrac{e^{-\eta\,z}}{4}(y^{2}-x^{2})$ (5.6)
@@
Again, the non deformed limit gives us back the non-deformed equations  (4.22). A symplectic realisation for this bi-Hamiltonian system is given by
$\begin{array}[]{l}x=\dfrac{1}{\eta}\sqrt{\dfrac{2-4\lambda}{2\lambda-1}}\sqrt{-(e^{\eta\,p}-1)^{2}\lambda-2\eta^{2}\,e^{\eta\,p}|k|}\sinh(\sqrt{1-2\lambda}\,q)\\ \\ y=\dfrac{2}{\eta\,\sqrt{2\lambda-1}}\sqrt{-e^{\eta\,p}(\eta^{2}\,|k|+\lambda(\cosh\eta\,p-1))}\cosh(\sqrt{1-2\lambda}\,q)\\ \\ z=p.\end{array}$ (5.9)
This realisation is valid for any value of $\lambda\neq 1/2$, that is, it is a family of parameterisations of the Poisson leaves of the deformed Poisson pencil (4.30).
Refer to captionRefer to caption
@@
It is interesting to comment on some aspects of the deformed dynamics of this system. Firstly, we note that although the deformed system (5.8) is symmetric under the change $x\leftrightarrow y$, this symmetry is not manifest in the numerical integration shown in Figure 3. This is due to the fact that the initial conditions for $x$ and $y$ have to be different. In fact, as it is easy to see from the fact that (5.6) is a constant of the motion, there cannot be closed orbits of the system such that $x(0)=y(0)$, thus preventing the existence of orbits reflecting the $x\leftrightarrow y$ symmetry of the system. Secondly, a closer look at the third equation $\dot{z}=-e^{-\eta\,z}xy$ shows that if $\eta$ is large and positive, then when $z$ becomes positive, the (closed) trajectories will reach a turning point since $\dot{z}\to 0$. A similar situation appears when $\eta$ is large and negative, and $z$ becomes negative. Both of these situations are clearly visible in the red and blue trajectories, respectively, from Figure 3.
@@ 6 Coupled systems and cluster variables: the AB case
In order to perform a detailed analysis of the $N=2$ coupled systems let us introduce for simplicity the following change of variables for the deformed AB case (4.30)
@@
Under the transformation (6.1) the deformed Poisson brackets reads
@@
Applying the transformation (6.1) to the AB deformed Casimir function (4.31) we arrive to the following expression
@@
Using the Poisson pencil (6.2) with $\lambda=1$, the equations in this coordinates are given by
@@
In order to compute the N=2 coupling we have to take into account the coproduct (4.4), which in the new variables (6.1) becomes
@@
From (6.7) we can define the so called cluster variables
@@
we compute the $N=2$ coupled system for the $\lambda=1$ case. Using the set of coordinates $\{x_{+},y_{+},z_{+},x_{1},y_{1},z_{1}\}$, the system of ODEs read
@@
Therefore we note that the dynamics for the cluster variables $\{x_{+},y_{+},z_{+}\}$ is exactly the same as the original system, while the dynamics for the other variables ($\{x_{1},y_{1},z_{1}\}$ in this case) is much more involved. It should be noted that this is a general feature of our formalism: cluster variables give a set of coordinates (canonically defined by means of the coproduct of the original variables) in which we recover a subsystem with identical dynamics to the original one. However, the coupled system is much more involved, and it is reflected in the dynamics for the rest of the coordinates we choose, which in general involves all of the variables.
A similar construction can be done for the $\lambda=0$ case. From (6.2) we have the following deformed Poisson-Lie brackets
@@
$\mathcal{H}_{0,\,\eta}\equiv-\mathcal{C}_{1,\eta}=\dfrac{4(\cosh(\eta\,z^{\prime})-1)+\eta^{2}\left((x^{\prime})^{2}+(y^{\prime})^{2}\right)}{4\,\eta^{2}}$ (6.12)
@@
Therefore, as we already know, the one-copy system is bi-Hamiltonian. This will clearly change when we consider the $N=2$ coupled system defined by the deformed coproduct (6.7) of the Hamiltonian function (6.12), namely
@@
$\begin{array}[]{ll}\begin{cases}\dot{x}_{+}=\dfrac{y_{+}\,\sinh(\eta\,z_{+})}{\eta}\\ \\ \dot{y}_{+}=\dfrac{x_{+}\,\sinh(\eta\,z_{+})}{\eta}\\ \\ \dot{z}_{+}=-x_{+}\,y_{+}\end{cases}\hskip 14.22636pt\begin{cases}\dot{x}^{\prime}_{1}=\dfrac{y^{\prime}_{1}}{4\,\eta}\left(4\,\sinh(\eta\,z_{+})+e^{\frac{\eta}{2}(z_{+}-z^{\prime}_{1})}\,\eta^{2}(x_{+}x^{\prime}_{1}+y_{+}y_{1})-\eta^{2}(x_{+}^{2}+y_{+}^{2})\right)\\ \\ \dot{y}^{\prime}_{1}=\dfrac{x_{1}}{4\,\eta}\left(4\,\sinh(\eta\,z_{+})+e^{\frac{\eta}{2}(z_{+}-z^{\prime}_{1})}\,\eta^{2}(x_{+}x^{\prime}_{1}+y_{+}y^{\prime}_{1})-\eta^{2}(x_{+}^{2}+y_{+}^{2})\right)\\ \\ \dot{z}^{\prime}_{1}=-\dfrac{1}{2}e^{\frac{\eta}{2}\,(z_{+}-z^{\prime}_{1})}(x_{+}y^{\prime}_{1}+y_{+}x^{\prime}_{1}).\end{cases}\end{array}$ (6.16)
As we can straightforwardly see, whilst the subsystem defined by $\{x_{+},y_{+},z_{+}\}$ is common for the two $\lambda$-cases (as it should be, since this is a direct consequence of the definition of the cluster variables), the equations for the variables $\{{x}_{1},{y}_{1},{z}_{1}\}$ is different for $\lambda=0$ and $\lambda=1$. Therefore, the coupling procedure defined by means of the Poisson-Lie group structure breaks the initial bi-Hamiltonian character of the system.

### v2

@@ 4.1.2 The Heisenberg-Weyl deformation
which is isomorphic to the Heisenberg-Weyl algebra. Following a similar procedure as in the previous case, we start with the faithful representation
@@
The compatible Poisson-Lie structure can be straightforwardly computed, and its fundamental Poisson brackets read
$\begin{array}[]{lll}\{x,y\}_{\eta}=0,&\{x,z\}_{\eta}=\eta\,x+y,&\{y,z\}_{\eta}=x+\eta\,y.\end{array}$ (4.13)
@@
Again, it is straightforward to check that in the limit $\eta\to 0$ we recover, up to a constant, the Poisson version of the $(1+1)$-dimensional Poincaré Lie algebra (3.5) and the Casimir function (2.6).
@@ 4.2 Bi-Hamiltonian deformations
Integrable deformations of Case B coming independently from each of its two Lie-Poisson Hamiltonian structures could be obtained by a completely analogous procedure to the one presented above, and will be based on the complete classification of Lie bialgebra structures for the centrally extended (1+1) Poincaré and $so(3)$ algebras. For the sake of brevity we shall not discuss them in this paper in detail. However, Rikitake system B is bi-Hamiltonian and thus the existence of the Poisson pencil (2.18) allows in principle to construct bi-Hamiltonian integrable deformations by following the prodedure introduced in [7, 8]. The essential point is that these deformations are possible if and only if there exist common cocycles $\delta$ for the two Lie algebras underlying the bi-Hamiltonian structure, namely the centrally extended Poincaré and $so(3)$ Lie algebras. In the following we prove that Case B does not admit bi-Hamiltonian integrable deformations since this condition cannot be fulfilled. Nevertheless, we will also show how Case AB does admit such a common cocycle, and we will compute and analyse a particular bi-Hamiltonian deformation for this system.
@@
We set a generic pre-cocommutator map $\delta:\mathfrak{g}\to\mathfrak{g}\wedge\mathfrak{g}$
$\begin{array}[]{l}\delta(X)=\eta\,(a_{1}\,X\wedge Y+a_{2}\,X\wedge Z+a_{3}\,X\wedge W+a_{4}\,Y\wedge Z+a_{5}\,Y\wedge W+a_{6}\,Z\wedge W),\\ \delta(Y)=\eta\,(b_{1}\,X\wedge Y+b_{2}\,X\wedge Z+b_{3}\,X\wedge W+b_{4}\,Y\wedge Z+b_{5}\,Y\wedge W+b_{6}\,Z\wedge W),\\ \delta(Z)=\eta\,(c_{1}\,X\wedge Y+c_{2}\,X\wedge Z+c_{3}\,X\wedge W+c_{4}\,Y\wedge Z+c_{5}\,Y\wedge W+c_{6}\,Z\wedge W),\\ \delta(W)=\eta\,(d_{1}\,X\wedge Y+d_{2}\,X\wedge Z+d_{3}\,X\wedge W+d_{4}\,Y\wedge Z+d_{5}\,Y\wedge W+d_{6}\,Z\wedge W),\\ \end{array}$ (4.17)
and we impose the cocycle condition
@@
$\begin{array}[]{l}\delta(X)=\eta\left((1-2\lambda)b_{3}\,Y\wedge W+b_{4}\,X\wedge Z-b_{4}\dfrac{\beta(1-\lambda)}{\lambda}X\wedge W+(2\beta(1-\lambda)c_{2}-2\lambda c_{3})Z\wedge W-c_{4}\,X\wedge Y\right),\\ \\ \delta(Y)=\eta\left(b_{3}\,X\wedge W+b_{4}\,Y\wedge Z-b_{4}\dfrac{\beta(\lambda-1)}{\lambda}Y\wedge W+b_{6}Z\wedge W+c_{2}X\wedge Y\right),\\ \\ \delta(Z)=\eta\left(c_{2}X\wedge Z+c_{3}X\wedge W+c_{4}Y\wedge Z+\left(\dfrac{(2\lambda-1)}{2\lambda}b_{6}+\dfrac{\beta(1-\lambda)}{\lambda}c_{4}\right)Y\wedge W\right),\\ \\ \delta(W)=0.\end{array}$ (4.19)
Since the cocommutator $\delta$ (4.19) should be defined for the complete Poisson pencil, including $\lambda=0$, we obtain that $b_{4}=0$, $b_{6}=0$, and $c_{4}=0$. Therefore we have
$\begin{array}[]{l}\delta(X)=\eta\left((1-2\lambda)b_{3}\,Y\wedge W+(2\beta(1-\lambda)c_{2}-2\lambda c_{3})Z\wedge W\right),\\ \\ \delta(Y)=\eta\left(b_{3}\,X\wedge W+c_{2}X\wedge Y\right),\\ \\ \delta(Z)=\eta\left(c_{2}X\wedge Z+c_{3}X\wedge W\right),\\ \\ \delta(W)=0.\end{array}$ (4.20)
The only remaining $\lambda$-independent solution is then obtained if $b_{3}=0$ and $c_{3}=-\beta c_{2}$,
@@
However, the co-Jacobi condition implies that the only cocommutator map is the trivial one $\delta(X)=\delta(Y)=\delta(Z)=\delta(W)=0$. Therefore, there are no non-trivial bi-Hamiltonian Poisson-Lie integrable deformations of the system (4.15).
@@ 4.2.1 Bi-Hamiltonian deformation of the Rikitake AB system
We have showed that Case A does not admit (non-trivial) bi-Hamiltonian deformations. However, when $\alpha=\mu=\beta=0$ (Case AB), the system (2.19) does admit such deformations. This case is given by the system of ODEs
@@
    First Hamiltonian structure: The (1+1)-Poincaré Lie-Poisson algebra
    where the Casimir function is given by $\mathcal{C}_{0}=\dfrac{1}{4}(y^{2}-x^{2})$, together with the Hamiltonian function
@@
    Second Hamiltonian structure: The $\mathfrak{so}(3)$ Lie-Poisson algebra
    $\begin{array}[]{lll}\{x,y\}_{1}=2z,&\{x,z\}_{1}=-y,&\{y,z\}_{1}=x,\end{array}$ (4.25)
@@
It is easy to see that, in contradistinction to Case A, now there exist $\lambda$-independent common cocycles, like for instance, the one given in (3.6),
@@
The ‘book’ Lie group can be embedded in $\mathrm{GL}(3,\mathbb{R})$ as shown in (4.2). Using this parametrisation, the coproduct $\Delta_{\eta}:\mathcal{C}^{\infty}(G^{\ast})\to\mathcal{C}^{\infty}(G^{\ast})\otimes\mathcal{C}^{\infty}(G^{\ast})$ is given by (4.4). A standard computation based on [27] shows that the unique Poisson structure on $G^{\ast}$ compatible with $\Delta_{\eta}$ with linearization given by (4.27) and therefore endowing $\mathcal{C}^{\infty}(G^{\ast})$ with a Poisson-Hopf algebra is given by the fundamental Poisson brackets
@@
The Casimir function for this Poisson-Lie group structure is given by
@@
As expected, when $\lambda=0$ and $\lambda=1$, we obtain Poisson-Lie deformations of the Lie-Poisson (1+1)-Poincaré (4.23) and $\mathfrak{so}(3)$ (4.25) Lie algebras, respectively:
@@
    Note that the only deformation in the Poisson brackets is produced for the $\{x,y\}_{0,\eta}$ bracket whose $\eta\rightarrow 0$ limit vanishes. In this case the deformed Casimir (4.31) reduces to
@@
    Moreover, the Casimir function is found to be
@@ 5.1 Case A
$\mathcal{H}^{A}=\dfrac{1}{2}(x^{2}+z^{2})-\alpha\,\log(x+y).$ (5.1)
The undeformed system (2.5) is obtained from the (1+1) Poincaré algebra (2.4), as shown in Section 2. This system could be seen as a dynamical system defined on an Abelian (trivial cocommutator) Poisson-Lie group. Therefore $\eta$-deformations, or equivalently, non-Abelian dual Poisson-Lie structures, define integrable deformations of this system. In this case, both Poisson-Lie structures explicitly given in the previous Section define the following deformed Rikitake systems.
$\bullet$ The “book” group deformation:
@@
A symplectic realisation of this deformed system corresponding to a value $\kappa$ of the Casimir function (4.7) is given by
@@
Note that in the $\eta\to 0$ limit we recover (2.7), both in the dynamical system and the symplectic realisation. Closed orbits of this deformed system are given in Figure 1.
@@
$\bullet$ The Heisenberg-Weyl deformation:
@@
$\begin{split}\dot{x}&=\{x,\mathcal{H}\}_{\eta}=z\,(y+\eta\,x),\\ \dot{y}&=\{y,\mathcal{H}\}_{\eta}=z\,(x+\eta\,y),\\ \dot{z}&=\{z,\mathcal{H}\}_{\eta}=\alpha-x\,y+\eta(\alpha-x^{2}).\end{split}$ (5.4)
A symplectic realisation of the Poisson algebra (4.13) with Casimir function (4.14) $\mathcal{C}_{\eta}=\kappa$ is given by
@@
The most significant difference between (5.3) and (5.5) is that in the former both $x$ and $y$ are functions of $q$ and $p$, while in the latter only functions of the $q$ variables appear.
@@
Closed trajectories for this system are presented in Figure 2. Note that trajectories with large positive values of the deformation parameter $\eta$ tend to be contained in a plane, while for large negative values of the deformation parameter $\eta$ their 3-dimensional nature is manifest.
@@ 5.2 Case AB
If we want to consider the bi-Hamiltonian deformation, we have to take as Hamiltonian either
$\mathcal{H}_{1,\eta}=\mathcal{C}_{0,\eta}=\dfrac{e^{-\eta\,z}}{4}(y^{2}-x^{2}),$ (5.6)
@@
Again, the non deformed limit leads to the non-deformed equations  (4.22). A symplectic realization for this bi-Hamiltonian system is given by
$\begin{array}[]{l}x=\dfrac{1}{\eta}\sqrt{\dfrac{2-4\lambda}{2\lambda-1}}\sqrt{-(e^{\eta\,p}-1)^{2}\lambda-2\eta^{2}\,e^{\eta\,p}|k|}\sinh(\sqrt{1-2\lambda}\,q),\\ \\ y=\dfrac{2}{\eta\,\sqrt{2\lambda-1}}\sqrt{-e^{\eta\,p}(\eta^{2}\,|k|+\lambda(\cosh\eta\,p-1))}\cosh(\sqrt{1-2\lambda}\,q),\\ \\ z=p,\end{array}$ (5.9)
provided that $\lambda\neq 1/2$. That is, (5.9) provides a family of parameterizations of the Poisson leaves of the deformed Poisson pencil (4.30).
@@
It is interesting to comment on some aspects of the deformed dynamics of this system. Firstly, we note that although the deformed system (5.8) is symmetric under the change $x\leftrightarrow y$, this symmetry is not manifest in the numerical integration shown in Figure 3. This is due to the fact that the initial conditions for $x$ and $y$ have to be different. In fact, as it is easy to see from the fact that (5.6) is a constant of the motion, there cannot exist closed orbits of the system such that $x(0)=y(0)$, thus preventing the existence of orbits reflecting the $x\leftrightarrow y$ symmetry of the system. Secondly, a closer look at the third equation $\dot{z}=-e^{-\eta\,z}xy$ shows that if $\eta$ is large and positive, then when $z$ becomes positive, the (closed) trajectories will reach a turning point since $\dot{z}\to 0$. A similar situation appears when $\eta$ is large and negative, and $z$ becomes negative. Both of these situations are clearly visible in the red and blue trajectories, respectively, from Figure 3.
@@ 6 Coupled systems and cluster variables: the AB case
In order to perform a detailed analysis of the coupled systems let us introduce for simplicity the following change of variables for the deformed AB case (4.30)
@@
Under the transformation (6.1) the deformed Poisson brackets read
@@
Applying the transformation (6.1) to the AB deformed Casimir function (4.31) we arrive at the following expression
@@
Using the Poisson pencil (6.2) with $\lambda=1$, the dynamical system in this coordinates is given by
@@
In order to compute the $N=2$ coupled system we have to take into account the coproduct (4.4), which in the new variables (6.1) becomes
@@
From (6.7) we can define the so called cluster variables [3] defined through the coproduct
@@
as the Hamiltonian for the $N=2$ coupled system in the $\lambda=1$ case, the set of coordinates $\{x_{+},y_{+},z_{+},x_{1},y_{1},z_{1}\}$ lead to the following system of coupled ODEs:
@@
Therefore we note that the dynamics for the cluster variables $\{x_{+},y_{+},z_{+}\}$ is exactly the same as the original system, while the dynamics for the other variables ($\{x_{1},y_{1},z_{1}\}$ in this case) is much more involved. It should be noted that this is a general feature of our formalism: cluster variables give a set of coordinates (canonically defined by means of the coproduct of the original variables) in which we recover a subsystem with identical dynamics to the original one. However, the “internal dynamics” of the coupled system is much more intricate, and this is reflected in the evolution for the rest of the coordinates we choose, which involves the six variables.
A similar construction can be performed for the $\lambda=0$ case. From (6.2) we have the following Poisson-Lie brackets
@@
$\mathcal{H}_{0,\,\eta}\equiv-\mathcal{C}_{1,\eta}=\dfrac{4(\cosh(\eta\,z^{\prime})-1)+\eta^{2}\left((x^{\prime})^{2}+(y^{\prime})^{2}\right)}{4\,\eta^{2}},$ (6.12)
@@
As we already know, the one-copy system is bi-Hamiltonian. This will clearly change when we consider the $N=2$ coupled system defined by the deformed coproduct (6.7) of the Hamiltonian function (6.12), namely
@@
$\begin{array}[]{ll}\begin{cases}\dot{x}_{+}=\dfrac{y_{+}\,\sinh(\eta\,z_{+})}{\eta},\\ \\ \dot{y}_{+}=\dfrac{x_{+}\,\sinh(\eta\,z_{+})}{\eta},\\ \\ \dot{z}_{+}=-x_{+}\,y_{+},\end{cases}\hskip 14.22636pt\begin{cases}\dot{x}^{\prime}_{1}=\dfrac{y^{\prime}_{1}}{4\,\eta}\left(4\,\sinh(\eta\,z_{+})+e^{\frac{\eta}{2}(z_{+}-z^{\prime}_{1})}\,\eta^{2}(x_{+}x^{\prime}_{1}+y_{+}y_{1})-\eta^{2}(x_{+}^{2}+y_{+}^{2})\right),\\ \\ \dot{y}^{\prime}_{1}=\dfrac{x_{1}}{4\,\eta}\left(4\,\sinh(\eta\,z_{+})+e^{\frac{\eta}{2}(z_{+}-z^{\prime}_{1})}\,\eta^{2}(x_{+}x^{\prime}_{1}+y_{+}y^{\prime}_{1})-\eta^{2}(x_{+}^{2}+y_{+}^{2})\right),\\ \\ \dot{z}^{\prime}_{1}=-\dfrac{1}{2}e^{\frac{\eta}{2}\,(z_{+}-z^{\prime}_{1})}(x_{+}y^{\prime}_{1}+y_{+}x^{\prime}_{1}).\end{cases}\end{array}$ (6.16)
As we can straightforwardly see, whilst the subsystem defined by $\{x_{+},y_{+},z_{+}\}$ is common for the two $\lambda$-cases (as it should be, since this is a direct consequence of the definition of the cluster variables), the equations for the variables $\{{x}_{1},{y}_{1},{z}_{1}\}$ are different for $\lambda=0$ and $\lambda=1$. Therefore, the coupling procedure defined by means of the Poisson-Lie group structure breaks the initial bi-Hamiltonian character of the system.